1. For parallel circuits, state the relationship between the voltage across the WHOLE circuit and the voltage across an INDIVIDUAL light bulb as more light bulbs are added.

2. For series circuits, state the relationship between the voltage across the WHOLE circuit and the voltage across an INDIVIDUAL light bulb as more light bulbs are added.
Answer #1

In Parallel circuits ,the voltage across whole circuit and across individual bulb be will be equal.In parallel connections voltage will be equal current would be divided between parallel branches. V=v1=v2=v3.

In Series  circuits ,the voltage across whole circuit will be equal to sum of  all bulbs in series  .Current through whole circuit and individual light bulbs will be equal . V=v1+v2+v3. v3=(V/(r1+r2+r3))* r3.